How would I factor the trinomial x^2+10x+24?
Answer by TimothyLamb(4379)

x^2 + 10x + 24 = 0
---
the above quadratic equation is in standard form, with a=1, b=10 and c=24
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 10 24
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
x = -4
x = -6
---
answer:
(x + 4)(x + 6)
---
you can also factor it by observation:
find two factors of 24 that when added together equal 10:
---
24 = 2*12 = 3*8 = 4*6
---
4 and 6 fit the bill: 4+6=10
---
answer by observation:
(x + 4)(x + 6)
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---